1. A firm is considering an investment in a new machine with a price of $15.6 million to replace its existing machine. The new machine is expected to have a four-year life. If the firm replaces the old machine with the new machine, it expects to save $6.3 million in operating costs each year over the next four years. The new machine will have no salvage value in four years. If the firm purchases the new machine, it will also need an investment of $250,000 in net working capital. The required return on the investment is 10 percent, and the tax rate is 39 percent. What are the NPV and IRR as a result of a decision to buy the new machine?
Answer #1

NPV=-15.6-250000/10^6+((6.3-15.6/4)*(1-39%)+15.6/4)/10%*(1-1/1.1^4)+250000/10^6*1/1.1^4=1.323911618 million

IRR:
-15.6-250000/10^6+((6.3-15.6/4)*(1-39%)+15.6/4)/IRR*(1-1/(1+IRR)^4)+250000/10^6*1/(1+IRR)^4=0

=>IRR=13.7771%